OpenStudy (anonymous):

is the limit of 3sqrt(x) -1 / sqrt(x) -1 as x approaches 1 equal to 3/2 ?

OpenStudy (anonymous):

I solved \[\frac{ \sqrt[3]{x} -1 }{ \sqrt{x} -1 }\]

OpenStudy (anonymous):

i solved the limit of this equation (sorry, forgot to add lim in front)

OpenStudy (anonymous):

so I solved it and got 3/2 is this correct?

OpenStudy (anonymous):

Yes, that is correct :) For future reference, if you ever want to check answers to questions like these, you can do that here: http://www.wolframalpha.com Just type, "limit of (x^(1/3) -1)/(sqrt(x) -1) as x tends to 1"! :-)

OpenStudy (anonymous):

thank you so much :)

OpenStudy (anonymous):

Actually, I misread your answer... you should have got \(\frac{2}{3}\). Is that what you got?

OpenStudy (anonymous):

hmm. I got 3/2

OpenStudy (anonymous):

here, I will type my procedure:

OpenStudy (anonymous):

\[\lim \frac{ x^3 -1 }{ x^2 -1 } = \lim \frac{ (x-1)(x^2+x+1) }{ (x-1)(x+1) } = \lim \frac{ x^2+x+1 }{ x+1}\]

OpenStudy (anonymous):

= 3/2

OpenStudy (anonymous):

Indeed, all of what you say is true :) ...but what makes you think that \[\lim_{x\rightarrow 1} \frac{\sqrt[3]{x}-1}{\sqrt{x}-1} \text{ and } \lim_{x\rightarrow 1} \frac{x^3-1}{x^2-1} \]are equal?

OpenStudy (anonymous):

i just took a wild guess. I'm not surprised I was wrong there then

OpenStudy (anonymous):

Wild guesses are extremely useful in maths, though it turns out that this time the limits are in fact, not equal. You had the right idea though. It is possible to factor the top (as a difference of two squares) and factor the bottom, and cancel, which should leave you with a fairly clear answer.

OpenStudy (anonymous):

hmm okay

OpenStudy (anonymous):

I don't know what else it would equal though..

OpenStudy (anonymous):

I think to do this question, you need to spot that the top is a difference of two squares: \[\sqrt[3]{x}-1 = (\sqrt[6]{x}-1)(\sqrt[6]{x}+1) \]and the bottom is the difference of two cubes. Do you know how to factorise the difference of two cubes?

OpenStudy (anonymous):

no, I did not learn this yet. we have never used 6root(x) ever :p

OpenStudy (anonymous):

Well, just because something hasn't explicitly been used doesn't mean it can't be :P If you want to find something that squares to a cube root, then that thing is the sixth root! If you don't want to use this method, have you by any chance come across l'Hopital's rule? I'm guessing not, but it's worth asking...

OpenStudy (anonymous):

no, haha. I have not come across that rule

OpenStudy (anonymous):

Hmm... it that case, I don't really see another way to this question. Are you familiar with indices? For example, that \(x^{\frac{1}{3}} = \sqrt[3]{x}\)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

And you are familiar with the laws of indices, such as \((x^a)^b=x^{(ab)}\)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

In that case, surely the statement \[(x^{\frac{1}{6}})^2 = x^{\frac{1}{3}}\]makes sense (all I'm doing is multiplying the powers)? You might not have seen 1/6 as a power before, but it still makes sense.

OpenStudy (anonymous):

okay ya, that makes sense

OpenStudy (anonymous):

Then all that says is \(\sqrt[3]{x} = (\sqrt[6]{x})^2\), so you can write \(\sqrt[3]{x}\) as a square. There you have a difference of two squares at the top.

OpenStudy (anonymous):

um, so I got lost here: \[\lim \frac{ x^{1/3} }{ (x-1)(x+1) }\]

OpenStudy (anonymous):

The bottom does not factor to (x+1)(x-1). \(\sqrt{x}-1\) and \(x^2-1\) are NOT the same thing! The second one does factor to (x+1)(x-1), but the first one does not. I think the key is probably to notice that \[\sqrt{x}-1 = x^{\frac{1}{2}} - 1 = (x^{\frac{1}{6}})^3 - (1)^3,\]which is the difference of two cubes.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

As for the top, you have \[\sqrt[3]{x} -1 = x^{\frac{1}{3}} - 1 = (x^{\frac{1}{6}})^2 - (1)^2,\]the difference of two squares.

OpenStudy (anonymous):

ok, yes what u are saying makes sense to me

OpenStudy (anonymous):

Good :-) So then \[\frac{\sqrt[3]{x} - 1}{\sqrt{x} - 1} = \frac{(x^{\frac{1}{6}}- 1)(x^{\frac{1}{6}}+ 1)}{(x^{\frac{1}{6}})^3- (1)^3}.\]I've factored the top since it's the difference of two squares. Now can you factor the difference of two cubes at the bottom?

OpenStudy (anonymous):

um, is it:

OpenStudy (anonymous):

\[(x^\frac{ 1 }{ 6 } -1)(x^\frac{ 1 }{ 3 } + x^ \frac{ 1 }{ 6 } +1)\]

OpenStudy (anonymous):

I have no idea :(

OpenStudy (anonymous):

Yup, that is correct :-)

OpenStudy (anonymous):

okay yay :)

OpenStudy (anonymous):

we can cross out the first parts of the numerator and denominator

OpenStudy (anonymous):

which is x^1/6 -1

OpenStudy (anonymous):

Exactly. Which leaves you with \[\frac{x^{\frac{1}{6}}+1}{x^{\frac{1}{3}}+x^{\frac{1}{6}}+1}\]and the limit of each of the terms in the top and the bottom is 1.

OpenStudy (anonymous):

Oh okay, and then that will give 2/3

OpenStudy (anonymous):

and the only reason why my initial procedure which gave 3/2 was wrong because my identities for the cube roots and square roots were not correct, right?

OpenStudy (anonymous):

I guess so. You incorrectly assumed that \[\frac{\sqrt[3]{x} -1}{\sqrt{x} -1} = \frac{x^3-1}{x^2-1}.\]

OpenStudy (anonymous):

yes, alright. thank you so much :)

OpenStudy (anonymous):

You're welcome :)